<p>We have recovered the beam at 01:49, after such a long time we had no target so we increased from 80K to 120K at 01:52. 02:01 update 120K-->95K as we have observed around that temperature always fluctuations and is safer.</p>
<p>Ring current 7.74e7 particles/s</p>
<p><strong>Merger, Tape server (WS3): </strong>1619 kB/s writing ok</p>
<p><strong>Leakage current (WS4): 3.26 uA </strong></p>
<p><strong>DAQ (WS2): temperature checked, clock status: ok</strong><strong>, adc calibration: ok, white rabbit check: failed</strong></p>
<p><strong>Online monitoring (WS 5): done</strong></p>
<p> </p> |